OpenStudy (anonymous):

Explain why (–4x)0 = 1, but –4x0 = –4.

OpenStudy (anonymous):

Because \[ (-4x)^0=((-4)^0)(x^0) \]And \[ -4x^0 = -4(x^0) \]

OpenStudy (anonymous):

Any nonzero number raised to zero is always equal to 1.